OpenStudy (anonymous):

Select the two values of x that are roots of the given polynomial below. 3x^2 + 8x - 3

OpenStudy (anonymous):

from what are we to select?

OpenStudy (anonymous):

first you factor it... it factors out to (3x-1)(x+3) then equal each one to 0 so.. 3x-1=0 1/3=x then do the next one so.. x+3=0 -3=x the roots would be x=-3 and x=1/3

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

your welcome